OpenStudy (anonymous):

Solve using the elimination method: 5/x +2/y=0 and 6/x+4/y=3

hero (hero):

I guess you didn't like my solution

OpenStudy (anonymous):

my hw website didnt accept it

hero (hero):

Oh, I see

hero (hero):

Next time, say something

hero (hero):

Allow me to try again

hero (hero):

Wait.... Did you find a value for y when you submitted your answer unto the website?

OpenStudy (anonymous):

the x value was wrong

hero (hero):

I see

hero (hero):

Try \[x = -\frac{4}{3}\]

OpenStudy (anonymous):

this prob is multiple choice and none of the answers are -4/3. with the other problem u solved for me, its fill in the blank. i put in the x value and it said it was incorrect.

hero (hero):

Well, sometimes the answer is in a different form. Answers can be expressed in more than one form. This is the reason why if you have a multiple choice problem, you should post the answer choices along with it. For the other problem, I suspect it is the same way. I suspect that the computer is looking for a different form of the same answer.

hero (hero):

I recommend that you post the answer choices so that I may see what form to put the answer in.

hero (hero):

I don't understand why you don't want to do this.

OpenStudy (anonymous):

answer is: x = -1 1/3, y = 8/15

hero (hero):

Yes, I know.... Guess what? -4/3 = -1 1/3

hero (hero):

I tried to explain to you that sometimes answers are in different form but the same

OpenStudy (anonymous):

im really tired right now so my brain is not functioning

hero (hero):

Next time, just post the answer choices and also say something like "the answer should be in the form of a mixed number"

OpenStudy (anonymous):

btw -4/3 is not the same as -1 1/3.... -2/3 is so you were wrong

OpenStudy (anonymous):

-1 1/3 is the same as -4/3, but is more confusing, for the exact reason that made you think it was -2/3.

OpenStudy (anonymous):

If it's a multiple choice question, you could just plug in all the answers until you find the one that works.